OpenStudy (anonymous):

Solve, using the substitution method. y-3x=7 21+9x=3y A.There is no solution B.There are an infinite number of solutions C.The solution is (1,10) D.The solution is (28,0)

OpenStudy (michele_laino):

hint: I solve the first equation for y, and I get: y=7+3x then I substitute that expression into the second equation, so we can write: \[21 + 9x = 3\left( {7 + 3x} \right)\] please solve that equation for x

OpenStudy (triciaal):

to use the substitution find the expression for one variable then use this expression for the variable in the second equation.you will have one equation with one unknown to solve for. when you have this value put it in the expression to get the other value.

OpenStudy (anonymous):

I think its C.

OpenStudy (anonymous):

then again Michele does make a point.

OpenStudy (triciaal):

what point? she did the same thing I outlined she did it for you. why not B?

OpenStudy (triciaal):

don't guess solve for x look at what @Michele_Laino wrote

OpenStudy (anonymous):

i think it has infinite solution...if u reaarage the equation, eq1: 3x-y+7=0 eq2: 9x-3y+21=0 which is equal to 3x-y+7=0 (taking 3 as a common factor) hence, we see that geometrically see that it represents the same line, hence there are infinite solution. in other words, we can say that lines represented by the equation are coincident, hence there infinte number of solutions